** SHOW ALL WORK !! **

.

Mary Family Farms produces watermelons. Their annual fixed cost to produce watermelons is $5,000. Their variable cost, including fertilizers and labor to produce one watermelon is $0.50. The Farm sells watermelons for $2.50 each.

.

a. If the Farm sells 20,000 watermelons annually, determine the following;

(i) the total annual cost

(ii) total annual revenue

(iii) Annual profit.

.

b. How many watermelons will the Farm need to sell in order to break even?  
Answer #1

a. i. Total annual cost

Total annual cost
Fixed annual cost 5000
Variable cost 0.5
Annual production 20000
10000
Total annual cost 15000

ii. Total annual revenue

selling cost 2.5
Annual sales 20000
Total annual revenue 2.5*20000 50000

iii. Annual profit

Annual profit
Total annual revunue - total annual cost 50000-15000 35000

b. Breakeven units

Fixed cost 5000
Selling price per unit 2.5
variable cost per unit 0.5
Difference 2.5-0.5 = 2
Break even units = fixed cost/ difference =5000/2 = 2500 units